If a 10-inch fish weighs 3 pounds, how much does a 15-inch fish weigh assuming the two fish are proportional? The answer is 10.125 pounds but I'm not sure how to get this.
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

If a 10-inch fish weighs 3 pounds, how much does a 15-inch fish weigh assuming the two fish are proportional? The answer is 10.125 pounds but I'm not sure how to get this.
---------------
Volume of a solid is proportional to the cube of its dimensions.
2x the radius ---> 8x the volume, etc.
---
A 15 inch fish is 1.5 x the length of a 10 inch fish.
1.5^3 = 3.375
3*3.375 = 10.125 pounds

Answer by stanbon(75887)

If a 10-inch fish weighs 3 pounds, how much does a 15-inch fish weigh assuming the two fish are proportional? The answer is 10.125 pounds but I'm not sure how to get this.
--------------------
wt/length = wt/length
x/15 = 3/10
------
x = 15*3/10
====
x = 45/10
---
x = 4.5 lbs
--------------------
Can't imagine where your 10.125 came from.
----------------
Cheers,
Stan H.
-----------------